OpenStudy (anonymous):

Use the first five terms of the exponential series and a calculator to approximate each value to the nearest hundreth: e^1.2

OpenStudy (anonymous):

cheat an use a calculator

OpenStudy (anonymous):

otherwise you need \[1+1.2+\frac{1.2^2}{2!}+\frac{1.2^3}{3!}+\frac{1.2^4}{4!}+\frac{1.2^5}{5!}\] which you will also use a calculator for